OpenStudy (anonymous):

Solve the system of equations using the elimination method and give your answer as an ordered pair: -2x + 2y = 2 -4y - 2x = -22

OpenStudy (anonymous):

multiply first equation by -1 and add - this ill eliminate x: 2x - 2y = -2 -2x - 4y = -22 add -6y = -24 y = 4 can you continue with this?